# RanDumb: A Simple Approach that Questions the Efficacy of Continual Representation Learning

### v1

@@ 3.2 RanDumb: Analysis
Ablating Components of RanDumb. We ablate the contribution of only using Random Fourier features for embedding and decorrelation to the overall performance of RanDumb in Table 9 (top). Ablating the decorrelation and relying solely on random Fourier features, colloquially dubbed Kernel-NCM, has performance drops ranging from 6-25% across the datasets. Replacing random Fourier features with raw features, colloquially the SLDA baseline, leads to pronounced drop in performance ranging from 3-14% across the datasets. Moreover, ablating both components, resulting in the base nearest class mean classifier exhibits the poorest performance with an average reduction of 17%. The poor performance when ablating both decorrelation and embedding highlights their interdependence.
Figure 3: Impact of Embedding Dimensions. We show the variation of accuracy of RanDumb with embedding dimensionality across datasets. We see decreasing marginal returns with increasing dimensions, obtaining relatively minor improvements after 15K dimensions.
@@
Impact of Embedding Dimensions. We vary the dimensions of the random Fourier features ranging from compressing 3K input dimensions to 1K to projecting it to 25K dimensions and evaluate its impact on performance in Figure 3. Surprisingly, the random projection to a 3x compressed 1K dimensional space allows for significant performance improvement over not using embedding, given in Table 9 (top). Furthermore, increasing the dimension from 1K to 25K results in improvements of 3.6%, 10.4%, 7.0%, and 2.5% on MNIST, CIFAR10, CIFAR100, and TinyImageNet respectively. Increasing the embedding sizes beyond 15K, however, only results in modest improvements of 0.1%, 1.4%, 1.1% and 0.2% on the same datasets, indicating 15K dimensions would be a good point for a performance-computational cost tradeoff.
@@
Impact of Flip Augmentation. We evaluate the impact of adding the flip augmentation on the performance of RanDumb in Table 9 (middle). Note that MNIST was not augmented. Augmentation provided large gains of 3.1% on CIFAR10, 1.7% on CIFAR100, and 0.4% on TinyImageNet. We did not try further augmenting the data with RandomCrop transform as done with standard augmentations.
@@
Impact of Varying Ridge Parameter. All prior experiments use a ridge parameter ($\lambda$) that increases with dataset complexity: $\lambda=10^{-6}$ for MNIST, $10^{-5}$ for CIFAR10 and CIFAR100, and $10^{-4}$ for TinyImageNet and miniImageNet. Table 9 (middle) shows the effect of varying $\lambda$ on RanDumb’s performance. With a smaller $\lambda=10^{-6}$, CIFAR10, CIFAR100, TinyImageNet and miniImageNet all exhibit minor drops of 0.1%-1.7%, 0.8%, 0.8%. Increasing shrinkage to a $\lambda=10^{-4}$ reduces CIFAR10 and CIFAR100 performance more substantially by 3% and 2.5% versus their optimal $\lambda=10^{-5}$. On the other hand, this larger $\lambda$ leads to improvements of 0.5% and 1.8% on TinyImageNet and miniImageNet. This aligns with the trend that datasets with greater complexity benefit from more regularisation, with the optimal $\lambda$ surfacing as a result of balancing under- and over-regularisation effects.
@@
Comparison with RanPAC. We compared our random Fourier features with projection based on a random weight matrix and ReLU [30] (RP+ReLU) in Table 9 (middle) with their best embedding size. Our method performs significantly better on each dataset, averaging a gain of 3.4%. We believe this is attributable to using a theoretical grounded random projection compared to RanPAC [30], which is specifically designed for the setting using pretrained transformers.
@@
Comparisons across Architectures. In table 7 (right), we compare whether using random Fourier features as embeddings outperforms models across various architectures for continual representation learning. We use experience replay (ER) baseline in the task-incremental CIFAR100 setup (for details, see Mirzadeh et al. [34] as it differs significantly from earlier setups). Our comparison involved across various architectures. The findings revealed that RanDumb surpassed the performance of nearly all considered architectures, and achieved close to 94% of the joint multi-task performance. This suggests that RanDumb outperforms continual representation learning across a wide range of architectures.
@@
Conclusion. Overall, both random embedding and decorrelation are critical components in the performance of RanDumb. Using random fourier features is substantially better than RanPAC. Lastly, one can substantially reduce the embedding dimension without a large drop in performance for large gains in computational cost and RAM usage, additional augmentation may further significantly help performance and optimal shrinkage parameter increases with dataset complexity. RanDumb outperforms continual representation learning across a wide range of architectures.
@@ 4 Discussion and Concluding Remarks
Our investigation reveals a surprising result, empirically demonstrating that currently the representation gain is negative, i.e. $RepGain(\mathcal{D})<0$ — simply using random embedding (RanDumb) consistently outperforms learned representations from methods specifically designed for continual training. Furthermore, using fixed pretrained networks consistently outperforms continually finetuned method. Using random/pretrained features also recovers 70-90% of the gap to joint learning, leaving limited room for improvement in representation learning techniques. Our investigation is both surprising and alarming as it questions our understanding of how to effectively design and train models that require efficient continual representation learning, and necessitates a re-investigation of the widely explored problem formulation itself. We believe adoption of computationally bounded scenarios without memory constraints [41, 40, 17] could be a promising way forward.

### v2

@@ 3.2 Analysis of RanDumb
Ablating Components of RanDumb. We ablate the contribution of only using Random Fourier features for embedding and decorrelation to the overall performance of RanDumb in Table 7 (left, top). Ablating the decorrelation and relying solely on random Fourier features, colloquially dubbed Kernel-NCM, has performance drops ranging from 6-25% across the datasets. Replacing random Fourier features with raw features, ie. the SLDA baseline, leads to pronounced drop in performance ranging from 3-14% across the datasets. Moreover, ablating both components results in the base nearest class mean classifier, and exhibits the poorest performance with an average reduction of 17%. Therefore, both decorrelation and random embedding are crucial for RanDumb.
@@
Impact of Embedding Dimensions. We vary the dimensions of the random Fourier features ranging from compressing 3K input dimensions to 1K to projecting it to 25K dimensions and evaluate its impact on performance in Figure 3. Surprisingly, the random projection to a 3x compressed 1K dimensional space allows for significant performance improvement over not using embedding, given in Table 7 (left, top). Furthermore, increasing the dimension from 1K to 25K results in improvements of 3.6%, 10.4%, 7.0%, and 2.5% on MNIST, CIFAR10, CIFAR100, and TinyImageNet respectively. Increasing the embedding sizes beyond 15K, however, only results in modest improvements of 0.1%, 1.4%, 1.1% and 0.2% on the same datasets, indicating 15K dimensions would be a good point for a performance-computational cost tradeoff.
@@
Impact of Flip Augmentation. We evaluate the impact of adding the flip augmentation on the performance of RanDumb in Table 7 (left, middle). Note that MNIST was not augmented. Augmentation provided large gains of 3.1% on CIFAR10, 1.7% on CIFAR100, and 0.4% on TinyImageNet. We did not augment the data further with RandomCrop transform as done with standard augmentations.
@@
Impact of Varying Ridge Parameter. All prior experiments use a ridge parameter ($\lambda$) that increases with dataset complexity: $\lambda=10^{-6}$ for MNIST, $10^{-5}$ for CIFAR10 and CIFAR100, and $10^{-4}$ for TinyImageNet and miniImageNet. Table 7 (left, middle) shows the effect of varying $\lambda$ on RanDumb’s performance. With a smaller $\lambda=10^{-6}$, CIFAR10, CIFAR100, TinyImageNet and miniImageNet all exhibit minor drops of 0.1%-1.7%, 0.8%, 0.8%. Increasing shrinkage to a $\lambda=10^{-4}$ reduces CIFAR10 and CIFAR100 performance more substantially by 3% and 2.5% versus their optimal $\lambda=10^{-5}$. On the other hand, this larger $\lambda$ leads to improvements of 0.5% and 1.8% on TinyImageNet and miniImageNet. This aligns with the trend that datasets with greater complexity benefit from more regularisation, with the optimal $\lambda$ balancing under- and over-regularisation effects.
Figure 3: Accuracy of RanDumb with respect to embedding dimensionality across datasets.
@@
Comparison with Extreme Learning Machines. We compared our random Fourier features with random projections based extreme learning machines, as recently adapted to continual learning by RP+ReLU [41] in Table 7 (left, bottom) with their best embedding size. Our method performs significantly better on each dataset, averaging a gain of 3.4%.
@@
Comparisons across Architectures. In table 7 (right), we compare whether using random Fourier features as embeddings outperforms models across various architectures for continual representation learning. We use experience replay (ER) baseline in the task-incremental CIFAR100 setup (for details, see Mirzadeh et al. [45] as it differs significantly from earlier setups). Our comparison spanned various architectures. The findings revealed that RanDumb surpassed the performance of nearly all considered architectures, and achieved close to 94% of the joint multi-task performance. This suggests that RanDumb outperforms continual representation learning across architectures.
@@
Conclusion. Overall, both random embedding and decorrelation are critical components in the performance of RanDumb. Using random Fourier features is substantially better than RanPAC. Lastly, one can substantially reduce the embedding dimension without a large drop in performance for large gains in computational cost, additional augmentation may further significantly help performance and optimal shrinkage parameter increases with dataset complexity. RanDumb outperforms continual representation learning across a wide range of architectures.
4 Related Work & Equivalent Formulations
Random Representations. There have been extensive theoretical and empirical investigations into random representations in machine learning, compressed sensing, and other fields, often utilizing extreme learning machines [56, 14, 21] (see [30, 29] for a survey). Other investigations include efficient kernel methods using Fourier features and Nyström approximations [52, 69], and extensions to efficiently parameterize linear classifiers [2]. They are also embedded into deep networks [17, 35, 72, 16]. We tailored the already successful random fourier representations [52] to the problem at hand and applied to the online continual learning problem for the first time.
Continual Representation Learning. There are various works focusing on continual representation learning itself [53, 20, 39, 28], but they address the problem of alleviating the stability-plasticity dilemma in high-exemplar and offline continual learning scenarios where models are trained until convergence. In comparison, we focus on online and low-examplar regime.
Representation Learning Free Methods in CL. Several works have developed the idea of using fixed pretrained networks after adapting on the first task across various settings [50, 41, 24]. Our work contributes to this growing evidence, however, we do not perform first-task adaptation [47], and propose OAS-shrinked SLDA as structurally simplest but highly accurate continual linear classifier without any extra bells-and-whistles. Moreover, we are the first work to introduce a representation learning free method with random features for continually learning from scratch.
Equivalent formulations to RanDumb. If the classes are equiprobable, which is the case for most datasets here, nearest class mean classifier with the Mahalanobis distance metric is equivalent to linear discriminant analysis (LDA) classifier [43]. Hence, one could say RanDumb is exactly equivalent to a Streaming LDA classifier with an approximate RBF Kernel. Alternatively, one could think of the decorrelation operation as explicitly decorrelating the features with ZCA whitening [7].
@@ 5 Discussion and Concluding Remarks
Our investigation reveals a surprising result — simply using random embedding (RanDumb) consistently outperforms learned representations from methods specifically designed for online continual training. Furthermore, using random/pretrained features also recovers 70-90% of the gap to joint learning, leaving limited room for improvement in representation learning techniques on standard benchmarks. Overall, our investigation questions our understanding of how to effectively design and train models that require efficient continual representation learning, and necessitates a re-investigation of the widely explored problem formulation itself. We believe adoption of computationally bounded scenarios without memory constraints and corresponding benchmarks [51, 50, 22] could be a promising way forward.
Limitations & Future Directions. We currently do not provide theory or justification for why training dynamics of continual learning algorithms fails to effectively learn good representations; doing so would provide deeper insights into continual learning algorithms. Moreover, our proposed method, RanDumb with random Fourier features is limited in scope towards low-exemplar scenarios and online-continual learning. Extending studies on representation learning to high-exemplar and offline continual learning scenarios might be exciting directions to investigate.